OpenStudy (anonymous):

How to derive a simple difference equation from a differential equation? for example, the discrete form of logistic equation is n(t)=a/(1+b*n(t-1)). Mine equation is much complicated than the logistic equation. Can someone tell me how to derive the discrete form of the logistic equation? ??? then start from there, I can work out my problem. Thank you!!

OpenStudy (anonymous):

whats your equation

OpenStudy (anonymous):

actually it is a set of equations: dn/dt=a*n-b*n*m/(c+n) dm/dt=d*m*a*n/(c+n)-e*m a,b,c,d,e are constants

OpenStudy (anonymous):

sorry this is new to me, i need to read up on it do you have a website etc

OpenStudy (anonymous):

I am just looking for the right material to read on the web. but no luck at this moment.

OpenStudy (anonymous):

You need to solve these with laplace transforms, also matlab can solve systems of ordinary differential equations. move all terms involving the independent variable and its derivatives on one side and all constant terms on the other side, You can then use the laplace transform approach to solve the problem.